# Primordial Black Hole Formation via Inverted Bubble Collapse

### v1

@@ II PBH formation in the inverted bubble collapse
Figure 1: PBH fraction in dark matter, $\mathrm{d}f_{\mathrm{PBH}}/\mathrm{d}\ln M$. The red and blue solid lines correspond to the parameter sets (a) and (b) in Table. 1, respectively. The red dashed and red dot-dashed lines adopt $\alpha=10$ and $v=0.5$ based on parameter set (a), respectively. For the blue dashed line, we set $T_{\mathrm{PT}}$ equal to $T_{c}$ in the parameter set (b) and tune $c$ so that it overlaps with the blue solid line for larger masses. The green dots represent the PBH fraction $f_{\rm PBH}$ in the singlet-extended model, with the dotted line showing its variation with the singlet triple coupling. The orange and light blue regions are obtained assuming the lensing event observed in the HSC M31 observation Niikura et al. (2019a) and OGLE Niikura et al. (2019b); Mroz et al. (2017) are due to PBHs, respectively. The purple region is the joint posterior for them Sugiyama et al. (2023). The gray lines represent the observational upper bounds on the PBH abundance by EROS Tisserand et al. (2007), HSC Niikura et al. (2019a), OGLE-III+OGLE-IV Mróz et al. (2024a), OGLE-IV high cadence survey Mróz et al. (2024b). The evaporation constraint is taken from PBH bounds Green and Kavanagh (2021); Kavanagh (2019)
III Model
In this section, we apply the inverted bubble collapse mechanism to the EWPT by introducing a real singlet scalar $S$ to the SM Higgs sector. The most general renormalizable scalar potential is given by
$V(\Phi,S)=$ $-m_{\Phi}^{2}|\Phi|^{2}+\lambda|\Phi|^{4}+\mu_{\Phi S}|\Phi|^{2}S+\lambda_{\Phi S}|\Phi|^{2}S^{2}$
$+t_{S}S-m^{2}_{S}S^{2}+\mu_{S}S^{3}+\lambda_{S}S^{4},$ (17)
where $\Phi$ denotes the Higgs doublet.
We choose the parameters in the scalar potential such that the phase transition proceeds in two stages. Initially, due to a local minimum along the $S$ direction, the $S$ field undergoes a first-order phase transition with bubble nucleation, where the $S$ field takes a larger value inside the bubbles than outside. Subsequently, the Higgs field in the bulk (outside the bubbles) undergoes EWPT at time $t=t_{\mathrm{PT}}$, before the completion of the first-order phase transition in the $S$ direction. While the vacuum state with a non-zero Higgs vacuum expectation value (VEV) in the bulk is not necessarily the true vacuum at $t_{\mathrm{PT}}$, it evolves into the true vacuum by $t=t_{c}$. At this point, the bubble interior, where the Higgs field remains in its symmetric phase, becomes a false vacuum, triggering the bubbles to collapse and form PBHs within a few Hubble times.
@@ III Model
When $t_{c}$ is significantly delayed beyond $t_{\mathrm{PT}}$, the resultant PBH mass tends to be larger for two reasons: the continued bubble expansion increases the false vacuum energy inside, and the collapse exhibits more run-away behavior, leading to a larger efficiency parameter $\epsilon$. Additionally, as $t_{c}$ is delayed, the physical size of bubbles asymptotically approaches the sound horizon size at $t=t_{c}$. This behavior occurs because no new bubbles form after $t_{\mathrm{PT}}$, and existing bubbles gradually approach the sound horizon. As a consequence, these similarly-sized bubbles collapse to form PBHs with highly monochromatic mass distribution, resulting in a sharp mass peak.
@@
The PBH abundance shown in Fig. 1 corresponds to the parameter choice $v=246.2\,{\rm GeV},\ m_{h}=125.09\,{\rm GeV},\ m_{H}=280\,{\rm GeV},\,\theta=0,\ \lambda_{S}=0.586,\ \lambda_{\Phi S}=0.5,\ \mu_{S}=[-246.025,-246.1]\,{\rm GeV},\ t_{S}=0,\ \mu_{\Phi S}=0$, where $v$ is the EW VEV, $m_{h}\ (m_{H})$ is the mass of SM-like (heavy) Higgs boson, and $\theta$ is the mixing angle for the CP-even Higgs bosons. Due to the extremely narrow mass spectrum in this model, we represent the PBH fraction $f_{\mathrm{PBH}}$ by green dots in the figure (using the same numerical scale as $\mathrm{d}f_{\rm PBH}/\mathrm{d}\ln M$ indicated on the vertical axis), with its variation with the singlet triple coupling shown by the dotted lines (see Appendix A for the detailed spectrum). We have used CosmoTransitions Wainwright (2012) to evaluate the tunneling rate and $\Delta V$ in the singlet extension model. For instance, we obtain $T_{\mathrm{PT}}\approx 148$ GeV and $T_{c}\approx 36$ GeV for the lower green dot in Fig. 1. The existence of such viable parameters in this minimal setup suggests that our PBH formation mechanism can be naturally realized in a wide class of models with multi-step phase transitions. Further details of the model are provided in Appendix A.
IV Discussion and Conclusions
In this Letter, we have proposed a novel PBH formation mechanism through inverted bubble collapse, where sparsely nucleated bubbles in an incomplete first-order phase transition become false vacuum regions after a bulk phase transition. These false vacuum bubbles collapse to form PBHs, allowing for precise prediction of PBH abundance through model parameters owing to the preserved spherical symmetry.
@@ IV Discussion and Conclusions
In our inverted bubble collapse mechanism, PBH abundance depends on both the rate and timing of bubble nucleation before the bulk phase transition, making it sensitive to the scalar potential shape. The scalar potential parameters must be set appropriately; the first-order phase transition should remain incomplete, as bubble collisions would break spherical symmetry and complicate PBH formation. If PBHs constitute a significant fraction of dark matter, this parameter constraint may be explained by anthropic selection; a completed first-order transition would trap our universe in a false vacuum, triggering old inflation and diluting baryons, while without a first-order transition, no PBH dark matter would form.
The final PBH mass depends on the detailed properties of the phase transitions. When applied to the EWPT, the PBH mass can reach up to ${\cal O}(10^{-6}\,\text{--}\,10^{-5})M_{\odot}$, corresponding to the horizon mass around the EW scale. Note that, as in the case studied in the previous section, it could take some time for the interior of the bubbles to become the false vacuum, which tends to increase the final PBH masses that exhibit a sharp mass distribution peak. This mass range of ${\cal O}(10^{-6}\,\text{--}\,10^{-5})M_{\odot}$ is particularly interesting as it could potentially explain the microlensing events observed in the OGLE and Subaru HSC data Mroz et al. (2017); Niikura et al. (2019a, b), and it will also be probed by future observations.
Smaller PBH masses are also possible, and PBHs generated by this mechanism could account for all dark matter. We have demonstrated that the inverted bubble collapse mechanism can be easily implemented in the EWPT with a real singlet scalar. A comprehensive exploration of the full parameter space remains an interesting direction for future work.
The PBH abundance depends on the properties of the bulk phase transition, with masses generally increasing if the transition is of first order. Interestingly, the first-order phase transition in the bulk will generate gravitational waves that could be probed by the future LISA experiment Arun et al. (2022). Further investigation of baryogenesis scenarios Kuzmin et al. (1985); Cohen et al. (1993) within the inverted bubble collapse framework is left for future investigation.
@@ Acknowledgments
We thank Katsuya Hashino for useful discussions. This work is supported by JSPS Core-to-Core Program (grant number: JPJSCCA20200002) (F.T.), JSPS KAKENHI Grant Numbers 20H01894 (F.T.), 20H05851 (F.T.), 23KJ0088 (K.M.), 24K17039 (K.M.), and 23KJ0086 (K.S.). This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
Appendix A Singlet Extension Model for EWPT
[…]
In this Appendix, we give a more detailed explanation of the singlet extension model given by Eq. (17). The component fields of $\Phi$ and $S$ are given by
$\Phi=\left[\begin{array}[]{c}G^{+}\\ \frac{1}{\sqrt{2}}\big{(}v+\phi+iG^{0}\big{)}\end{array}\right],\quad S=v_{S}+s,$ (20)
where $v\simeq 246$ GeV and $v_{S}$ are the VEVs of $\Phi$ and $S$, respectively, and $G^{+}$ and $G^{0}$ are Nambu-Goldstone bosons that are absorbed by the weak gauge bosons. The shift of $S$, $S\to S+v^{\prime}_{S}$, does not change physics Kanemura et al. (2017). Due to this shift invariance, we take $v_{S}=0$ in the following discussion. The CP-even scalar states $\phi$ and $s$ mix with each other. The mass matrix in the weak basis $(s,\phi)$ is given by
${M}^{2}$ $=\left(\begin{array}[]{cc}{M}_{ss}^{2}&{M}_{sh}^{2}\\ {M}_{sh}^{2}&{M}_{hh}^{2}\\ \end{array}\right),$ (23)
$M_{ss}^{2}$ $=2m_{S}^{2}+\lambda_{\Phi S}v^{2},M_{sh}^{2}=\mu_{\Phi S}v,M_{hh}^{2}=2\lambda v^{2}.$ (24)

### v3

@@ II.4 PBH mass function
Figure 2: PBH fraction in dark matter, $\mathrm{d}f_{\mathrm{PBH}}/\mathrm{d}\ln M$. The red and blue solid lines correspond to the parameter sets (a) and (b) in Table. 1, respectively. The red dashed and red dot-dashed lines correspond to modifications of parameter set (a), where we set $\alpha=10$ and $v=0.5$, respectively. For the blue dashed line, we set $T_{\mathrm{PT}}$ equal to $T_{c}$ in the parameter set (b) and tune the parameter $c$ so that it overlaps with the blue solid line for larger masses. The green dots represent the PBH fraction $f_{\rm PBH}$ in the singlet-extended model, with the dotted line showing its variation with the triple coupling of $S_{1}$. The orange and light blue regions are obtained assuming the lensing event observed in the HSC M31 observation [60] and OGLE [61, 59] are due to PBHs, respectively. The purple region is the joint posterior for them [75]. The gray lines represent the observational upper bounds on the PBH abundance by EROS [76], HSC [60], OGLE-III+OGLE-IV [77], OGLE-IV high cadence survey [78]. The evaporation constraint is taken from PBH bounds [11, 79].
III Singlet Extended Model
III.1 Description of the model
In this section, we apply the IBC mechanism to the singlet extension of the SM where we introduce two real singlets $S_{1}$ and $S_{2}$. The scalar potential is given by
$V(\Phi,S_{1},S_{2})=$ $-m_{\Phi}^{2}|\Phi|^{2}+\lambda|\Phi|^{4}+\mu_{\Phi 1}|\Phi|^{2}S_{1}+\lambda_{\Phi 1}|\Phi|^{2}S^{2}_{1}+t_{1}S_{1}-m^{2}_{1}S^{2}_{1}+\mu_{1}S^{3}_{1}$
$+\lambda_{1}S^{4}_{1}-m_{2}^{2}S_{2}^{2}+\lambda_{2}S_{2}^{4}+\lambda_{\Phi 2}|\Phi|^{2}S_{2}^{2}+\lambda_{12}S_{1}^{2}S_{2}^{2}+\mu_{12}S_{1}S_{2}^{2},$ (16)
where $\Phi$ denotes the Higgs doublet, and the potential respects the $Z_{2}$ symmetry acting on $S_{2}$. [^4] The collider phenomenology in the extension of the SM with two real singlet scalars with a $Z_{2}\times Z_{2}^{\prime}$ discrete symmetry are studied in Ref. [80]. The component fields of $\Phi$ and $S_{i}$ are given by
$\Phi$ $=\left[\begin{array}[]{c}G^{+}\\ \frac{1}{\sqrt{2}}\big{(}v+\phi+iG^{0}\big{)}\end{array}\right]\;,\quad S_{i}=v_{i}+s_{i}\ (i=1,2)\;,$ (19)
where $v\simeq 246$ GeV and $v_{i}$ are the VEVs of $\Phi$ and $S_{i}$, respectively, and $G^{+}$ and $G^{0}$ are Nambu-Goldstone (NG) bosons that are absorbed by the weak gauge bosons. The shift of $S_{1}$, $S_{1}\to S_{1}+v^{\prime}_{1}$, does not change physics [81]. Using this shift transformation, we take $v_{1}=0$ in the following discussion. In the mass basis, three CP-even Higgs bosons are identified by the SM-like Higgs boson $(h)$ and heavy Higgs bosons ($H_{1}$ and $H_{2}$). By setting the tadpole conditions of the scalar potential (16) and diagonalizing the mass matrix, the original potential parameters can be expressed in terms of physical parameters, i.e., the masses of the Higgs bosons ($m_{h}$, $m_{H_{1}}$ and $m_{H_{2}}$) and the mixing angles $\theta_{i}$ $(i=1,2,3)$. The diagonalization of the mass matrix is presented in Appendix A. We choose the following twelve parameters as input parameters of the model:
$v\;,\quad m_{h}\;,\quad v_{2}\;,\quad m_{H_{1}}\;,\quad m_{H_{2}}\;,\quad\theta_{i}\;\ \ (i=1,2,3),\quad\mu_{1}\;,\quad\lambda_{1}\;,\quad\lambda_{\Phi 1}\;,\quad\lambda_{12}\;,\quad$ (20)
where the EW VEV and the SM-like Higgs boson mass are fixed to be $v=246.2\,{\rm GeV}$ and $m_{h}=125.09\,{\rm GeV}$.
III.2 Construction of effective potential
From the potential (16), one can calculate the effective potential, which consists of the three contributions,
$V_{\rm eff}(\phi,s_{i},T)=V_{0}(\phi,s_{i})+V_{\rm CW}(\phi,s_{i})+V_{T}(\phi,s_{i},T)~{}(i=1,2)\;,$ (21)
where $V_{0}(\phi,s_{i})$ is the tree-level potential, $V_{\rm CW}(\phi,s_{i})$ is Coleman-Weinberg potential up to the one-loop level [82], and $V_{T}(\phi,s_{i},T)$ is the finite-temperature effect of the effective potential [83]. The tree-level contribution $V_{0}$ can be straightforwardly obtained from the potential (16) with the replacement of $(\Phi,S_{i})\to(\phi,s_{i})$. The remaining parts $V_{\rm CW}$ and $V_{T}$ depend on the field-dependent masses $\bar{m}^{2}_{k}(\phi,s_{i})$ of the model. In terms of the field-dependent masses, $V_{\rm CW}$ and $V_{T}$ are given by
$V_{\text{CW}}$ $=\sum_{k}(-1)^{F_{k}}\frac{n_{k}}{64\pi^{2}}\bar{m}_{k}^{4}(\phi,s_{i})\left(\log{\frac{\bar{m}_{k}^{2}(\phi,s_{i})}{\mu^{2}_{R}}}-c_{k}\right)\,,$ (22)
$V_{T}$ $=\frac{T^{4}}{2\pi^{2}}\sum_{k}n_{k}J_{\mp}\left(\frac{\bar{m}_{k}(\phi,s_{i})}{T}\right)\,,$ (23)
where $n_{k}$ is the number of degrees of freedom of the particle species labeled by $k$, and the constant $c_{k}$ is 3/2 (5/6) for scalar bosons and fermions (gauge bosons). Field-dependent masses $\bar{m}_{k}(\phi,s_{i})$ are presented in Appendix B. The index $F_{k}$ is $0\ (1)$ for bosons (fermions). The function $J_{\mp}(x)$ is given by
[…]
with the minus (plus) sign being for the bosons (fermions). Renormalization of the Coleman-Weinberg potential is performed in the $\overline{\rm MS}$ scheme, by which the renormalization scale $\mu_{R}$ is introduced in the effective potential. [^5] The Coleman-Weinberg potential can also be evaluated by applying another renormalization scheme, such as on shell-like schemes. The renormalization scheme difference of the effective potential is discussed in Refs. [84, 85]. In our analysis, we set $\mu_{R}=m_{t}$.
We implement the thermal effects of the field-dependent masses by the replacement $\bar{m}^{2}_{k}(\phi,s_{i})\to\bar{M}^{2}_{k}(\phi,s_{i},T)$ in the effective potential $V_{\rm eff}$ [86]. The thermally corrected masses $\bar{M}^{2}_{k}(\phi,s_{i},T)$ involve the temperature dependent self-energy $\Pi_{k}(T)$, and those for the Higgs field and singlet scalar fields are given by
$\Pi_{\Phi}(T)$ $=T^{2}\Bigg{[}\frac{1}{16}\left(3g^{2}+g^{\prime 2}\right)+\frac{y_{t}^{2}}{4}+\frac{1}{6}\left(\frac{3\lambda}{2}+\frac{\lambda_{\Phi 1}}{2}+\frac{\lambda_{\Phi 2}}{2}\right)+\frac{\lambda}{4}\Bigg{]}\;,$ (25)
$\Pi_{S_{1}}(T)$ $=T^{2}\left[\lambda_{1}+\frac{1}{6}(\lambda_{12}+2\lambda_{\Phi 1})\right]\;,$ (26)
$\Pi_{S_{2}}(T)$ $=T^{2}\left[\frac{1}{6}(\lambda_{12}+2\lambda_{\Phi 2})+\lambda_{2}\right]\;.$ (27)
One then obtains a field-dependent mass matrix with the thermal corrections in the basis of $(\phi,s_{1},s_{2})$: $\bar{{M}}^{2}=\bar{{\cal M}}^{2}+{\rm diag}(\Pi_{\Phi},\Pi_{S_{1}},\Pi_{S_{2}})$, where ¯$\bar{{\cal M}}^{2}$ is the field-dependent mass matrix without thermal corrections in the basis of $(\phi,s_{1},s_{2})$ given in Appendix B. Diagonalizing $\bar{M}^{2}$ yields the thermally corrected masses of the CP-even Higgs bosons. The thermally corrected masses of the NG boson are obtained by $\bar{M}^{2}_{G^{0},G^{\pm}}=\bar{m}_{G^{0},G^{\pm}}^{2}+\Pi_{\Phi}$. The masses of the longitudinal mode of the weak gauge bosons are also replaced with thermally corrected ones [87].
III.3 PBH formation
Figure 3: Temperature dependence of three different minima. The blue (green) line shows the local minima appearing in the $S_{1}$ ($S_{2}$)-direction ($S_{1}$ ($S_{2}$)-vacuum). The orange line shows the minimum that exists at high temperatures.
[…]
Figure 5: The highly monochromatic PBH mass spectrum in the singlet-extension model. The spectra from top to bottom are calculated with different values of the triple coupling of $S_{1}$, $\mu_{1}=-190.5,-190.49,-190.48,-190.47,-190.46$ GeV, while other parameters are fixed as (III.3). Note that the PBH fraction shown by the green dots is about two orders of magnitude smaller than the peak height due to the extremely narrow width.
In our IBC mechanism, the PBH formation is related to phase transitions in the directions of $S_{1}$ and $S_{2}$ that occur before the EWPT. Therefore, we analyze the scalar potential by setting the CP-even component of $\Phi$ to be $\phi=0$. While the required delayed EWSB can be realized for a certain choice of the model parameters [67, 70], we focus on the phase transition dynamics in the singlet sector for our purpose.
To realize the IBC mechanism, we set the input model parameters as follows,
$v=246.2\,{\rm GeV},\ m_{h}=125.09\,{\rm GeV},\ m_{H_{1}}=223.1\,{\rm GeV},\,m_{H_{2}}=246\,{\rm GeV},\,v_{2}=70{\rm GeV},\,$
$\theta_{1}=\theta_{2}=\theta_{3}=0,\ \lambda_{1}=0.335,\ \lambda_{12}=0.5,\ \lambda_{\Phi 1}=-0.14,\ \mu_{1}=[-190.49,-190.48]\,{\rm GeV}\;.$ (28)
For these parameter choices, we have checked that the scalar potential satisfies the releavnt theoretical conditions, i.e., perturbative unitarity [80] and the stability condition of the potential [80]. Experimental constraints such as direct searches of the heavy Higgs boson and Higgs coupling measurements for the model with a single real singlet scalar field are discussed using the Run II data of LHC in the Appendix of Ref. [71]. The above parameter points satisfy all these experimental constraints.
[…]
@@ III.3 PBH formation
When $t_{c}$ is significantly delayed beyond $t_{\mathrm{PT}}$, the resultant PBH mass tends to be larger for two reasons: the continued bubble expansion increases the net false vacuum energy inside, and the collapse exhibits more runaway behavior, leading to a larger efficiency parameter $\epsilon$. Additionally, as $t_{c}$ is delayed, the physical size of bubbles asymptotically approaches the sound horizon size at $t=t_{c}$. This behavior occurs because no new bubbles form after $t_{\mathrm{PT}}$, and existing bubbles gradually approach the sound horizon. As a consequence, these similarly-sized bubbles collapse to form PBHs with highly monochromatic mass distribution, resulting in a sharp mass peak.
@@
The PBH abundance shown in Fig. 2 (represented by green dots and dotted line) corresponds to the parameter set (III.3). Due to the extremely narrow mass spectrum in this model, we represent the PBH fraction $f_{\mathrm{PBH}}$ by green dots in the figure (using the same numerical scale as $\mathrm{d}f_{\rm PBH}/\mathrm{d}\ln M$ indicated on the vertical axis), with its variation with the singlet triple coupling shown by the dotted lines as we will explain below. The existence of such viable parameters in this simple setup suggests that our PBH formation mechanism can be naturally realized in a wide class of models with multi-step phase transitions.
Fig. 5 shows the highly monochromatic PBH mass spectrum predicted in the singlet-extension model. It should be noted that due to the extremely narrow peak width, the PBH fraction in dark matter obtained by integrating the distribution (green dots) is about two orders of magnitude smaller than the peak height of the mass distribution. This spectrum is obtained by substituting the temperature-dependent bubble nucleation rate and $\Delta V$ calculated using CosmoTransition into Eq. (15). Here, the spectra from top to bottom correspond to different values of the triple coupling of $S_{1}$, $\mu_{1}=-190.5,-190.49,-190.48,-190.47,-190.46$ GeV, while keeping the other parameters fixed as in Eq. (III.3).
IV Discussion
@@
For successful PBH formation in our IBC scenario, the scalar potential parameters must be set appropriately; the FOPT should remain incomplete, as bubble collisions would break spherical symmetry and complicate PBH formation. If PBHs constitute a significant fraction of dark matter, this parameter constraint may be explained by anthropic selection; a completed FOPT would trap our universe in a false vacuum, triggering old inflation and diluting baryons, while without a FOPT, no PBH dark matter would form.
One of the key elements of our IBC mechanism is the runaway behavior of the collapsing bubbles. There is currently ongoing debate about establishing a definite criterion for the runaway bubble dynamics, but the bubbles in the singlet sector are likely the ones that exhibit the runaway behavior. Even if the bubble dynamics does not exhibit this runaway behavior, we may have successful PBH formation by introducing a short period of inflation [89, 90, 91] after the bubble nucleation. Then, the overdense regions in the interior of the bubbles are stretched beyond the horizon, and they could collapse into PBHs when the bubbles reenter the horizon, if the density perturbations are sizable. Note that the late-time inflation does not necessarily have to end inside bubbles, in which case the PBHs can still form when the size of the bubbles becomes smaller than the Hubble radius of the universe outside the bubbles. This is conceptually related to the PBH formation using vacuum bubbles [92, 93] (see also Ref. [94]).
Furthermore, while this work focuses on pre-EWPT phase transitions in the singlet sector, similar IBC mechanisms might operate during or after the EWPT in singlet or hidden sectors, potentially triggered by the EWPT itself. These scenarios offer additional pathways for PBH formation worthy of future investigation. Also, it is an interesting question if the baryogenesis can be incorporated in our example of the singlet-extended model, especially in a case where the bulk phase transition or the EWPT is of first order. Investigation of baryogenesis scenarios [95, 96] within this framework is left for future investigation.
@@ Acknowledgments
We thank Katsuya Hashino for useful discussions, and Wen-Yuan Ai for informing us of Ref. [72]. This work is supported by JSPS Core-to-Core Program (grant number: JPJSCCA20200002) (F.T.), JSPS KAKENHI Grant Numbers 20H01894 (F.T.), 20H05851 (F.T.), 23KJ0088 (K.M.), 24K17039 (K.M.), and 23KJ0086 (K.S.). This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
Appendix A Diagonalization of the mass matrix for the CP-even Higgs boson
In this Appendix, we perform the diagonalization of the mass matrix for the CP-even Higgs boson to obtain the relation between the original scalar potential parameter and physical parameters. The CP-even scalar states $\phi$ and $s_{i}$ mix with each other. The mass matrix in the weak basis $(\phi,s_{1},s_{2})$ is given by
${M}^{2}$ $=\left(\begin{array}[]{ccc}{M}_{\phi\phi}^{2}&{M}_{\phi 1}^{2}&{M}_{\phi 2}^{2}\\ {M}_{\phi 1}^{2}&{M}_{11}^{2}&{M}_{12}^{2}\\ {M}_{\phi 2}^{2}&{M}_{12}^{2}&{M}_{22}^{2}\\ \end{array}\right),$ (32)
$M_{\phi\phi}^{2}$ $=2\lambda v^{2},\ \ \ M_{\phi 1}^{2}=\mu_{\Phi 1}v,\ \ \ M_{\phi 2}^{2}=2\lambda_{\Phi 2}vv_{2},$ (33)
$M_{11}^{2}$ $=-2m_{S}^{2}+\lambda_{\Phi 1}v^{2}+2\lambda_{12}v_{2}^{2},\ \ \ M_{12}^{2}=2\mu_{12}v_{2},\ \ \ M_{22}^{2}=8\lambda_{2}v_{2}^{2},$ (34)